# Let the Experts Speak: Improving Survival Prediction & Calibration via Mixture-of-Experts Heads

### v1

@@ Abstract
Deep mixture-of-experts models have attracted a lot of attention for survival analysis problems, particularly for their ability to cluster similar patients together. In practice, grouping often comes at the expense of key metrics such calibration error and predictive accuracy. This is due to the restrictive inductive bias that mixture-of-experts imposes, that predictions for individual patients must look like predictions for the group they’re assigned to. Might we be able to discover patient group structure, where it exists, while improving calibration and predictive accuracy? In this work, we introduce several discrete-time deep mixture-of-experts (MoE) based architectures for survival analysis problems, one of which achieves all desiderata: clustering, calibration, and predictive accuracy. We show that a key differentiator between this array of MoEs is how expressive their experts are. We find that more expressive experts that tailor predictions per patient outperform experts that rely on fixed group prototypes.
@@ 1 Introduction
The first architecture, Fixed MoE, uses several experts where each expert learns an associated event distribution, which is then fixed across all patients. The second architecture, Adjustable MoE, again learns a prototypical event distribution per expert, but can be adjusted to the individual patient. The third architecture, Personalized MoE, uses experts that each form a custom event distribution for individual patients. Only the third architecture is able to cluster patients and outperform strong baseline models on calibration, concordance index, and time-dependent Brier score. The contributions of our work are as follows: (i) we introduce three discrete-time deep MoE based survival architectures, one of which achieves excellent clustering, calibration error, concordance index, and time-dependent Brier score and (ii) we report that the expressiveness of the experts is a key differentiator between discrete-time deep MoE-based survival analysis models, supported by a targeted set of experiments on these three architectures.
@@ 2 Related Work
Building survival models with flexible model classes, such as Gaussian processes (Alaa and van der Schaar, 2017) or neural networks (Ranganath et al., 2016), and rich distribution families (Miscouridou et al., 2018), yields better predictions of risk than standard linear Cox or accelerated failure-time (AFT) models. One can retain the structure of AFT models but improve prediction by using deep representations of the covariates (Zhong et al., 2021; Norman et al., 2024). Further generalizations that retain structure were introduced as conditional transformation models (Hothorn et al., 2018; Campanella et al., 2022) These choices improve expressivity to allow for better approximations of complicated survival distributions but do not reveal any structure in the data.
@@ 3 Methods
We now describe our three proposed deep mixture-of-experts (MoE) based survival architectures. All models are feedforward deep learning models that use information from a patient’s record (e.g., demographic data, physiological data, etc.) to forecast when they are likely to have a clinical event. Raw patient records $\mathbf{x}_{0}^{i}\in\mathbb{R}^{a}$ for patients $i=1\ldots N$ and $a$ the number of raw features are composed of categorical indicators (e.g., gender, etc.), and standardized continuous features (e.g., heart rate), meaning that for continuous features we subtract the mean and divide by the standard deviation. We learn embeddings for categorical indicators. After embedding lookups are concatenated with continuous features, this results in a $d$-dimensional feature vector that is fed to the model. Let $\mathbf{x}=\mathbf{x}^{i}_{\ell-1}\in\mathbb{R}^{h}$ be shorthand for the $h$-dimensional penultimate hidden state representation of our feedforward model with $\ell$ layers for the $i^{\text{th}}$ patient. We will now describe the $\ell^{\text{th}}$ layer (i.e., the final layer), which is the MoE head in all architectures. All methods are trained using a discrete-time Multitask Logistic Regression (MTLR) style loss function (Yu et al., 2011; Fotso, 2018) that predicts a monotone label sequence (“once event occurs, it stays on”). This loss function has been shown to be well-calibrated (Haider et al., 2020). The loss functions are described in Appendix Section B.
@@ 3.3 Personalized Mixture-of-Experts
The Personalized MoE architecture belongs to a collection of models that provide the most flexibility to the experts to form custom event distributions per patient (e.g., Deep Survival Machines (Nagpal et al., 2021a)). Since this architecture generates new expert distributions for each patient, we first project the final hidden state representaion $\mathbf{x}$ to a router representation with $\mathbf{x}_{r}=\mathbf{x}W^{\intercal}_{r}$ for $W_{r}\in\mathbb{R}^{h\times h}$ and an expert representation with $\mathbf{x}_{e}=\mathbf{x}W^{\intercal}_{e}$ for $W_{e}\in\mathbb{R}^{h\times h}$. We then divide the expert representation into $n$ evenly-sized chunks $\mathbf{x}_{e,k}$ for $k=1,\ldots,n$, which are each fed to a linear layer denoted $L_{k}\in\mathbb{R}^{m\times(h/n)}$ to form an event-distribution to obtain $M_{k}(\mathbf{x}_{e,k})=\mathbf{x}_{e,k}L_{k}^{\intercal}$, which collectively form the dynamic matrix of unnormalized densities over event times $M(\mathbf{x}_{e})\in\mathbb{R}^{n\times m}$. The final PMF is then
@@ 4 Experiments
We experiment with 3 datasets to probe the properties and capabilities of our proposed methods. Survival MNIST is a synthetic dataset (see Figure 2), which allows us to probe the models’ abilities to predict event times and recover latent groups. We censor 15% of examples. SUPPORT2 is a survival analysis dataset with 9,105 examples and $\sim$32% censoring (Connors et al., 1995). Sepsis is a larger dataset with 40,336 patient records and only 2,932 positive instances of sepsis, making this a very challenging anomaly detection survival analysis task (Reyna et al., 2020). For the Sepsis dataset, the first 100 hours of each patient’s ICU stay was summarized to perform a retrospective prediction, helpful when using partially labeled historical datasets to accelerate human labeling (Singer et al., 2016; Henry et al., 2019). Patients in the Sepsis dataset were administratively censored after 100 hours.
@@ 5 Results
We also analyze the clustering behavior of a Personalized MoE model trained on SUPPORT2 in Figure 5 to better understand how patient groups are formed on a real-world dataset. We see a variety of clusters sizes (Figure 5), indicating that Cluster 1 could possibly be merged into another cluster given its small size. We also find that patient risks are separated by cluster as evidenced by Figure 5. We then investigate kernel density estimates of continuous features (e.g., age in Figure 5) broken down by cluster. For categorical features, we define a $z$-score based on the proportion of patients in a cluster that have a given categorical feature value relative to the overall proportion of patients with that feature value (see Appendix Section J for more details). Importantly, these analyses allowed us to define succinct, clinician-informed signatures for each cluster, which are summarized in Table 2 and described in more detail in Appendix Section I. This indicates that the Personalized MoE model is able to discover clinically meaningful patient groups in real-world datasets.
@@ 6 Conclusion
We introduce three discrete-time deep mixture-of-experts (MoE) based survival architectures, one of which achieves excellent clustering, calibration error, and Brier scores at key points in time. We have shown that the expressiveness of the experts is a key differentiator between deep MoE-based survival analysis models, supported by a targeted set of experiments on these three architectures. We observe that the Personalized MoE model is able to deliver on all desiderata: clustering, calibration, and predictive accuracy as evidenced by results on three survival analysis datasets, including two real-world datasets. We explore sensitivity to the number of experts and find that the Personalized MoE model is the least sensitive to the number of experts, making it a robust choice for survival analysis tasks. We also analyze the routing behavior of the MoE models and use a synthetic survival analysis dataset to validate that the models are able to recover latent groups in the data. In practice, we find that the Personalized MoE model is able to discover clinically meaningful patient groups in real-world datasets. Overall, the Personalized MoE model provides a powerful and flexible approach to survival analysis that can adapt to the complexities of real-world clinical data and patient heterogeneity. It maintains strong performance across multiple metrics of interest to clinicians and researchers alike, including calibration and predictive accuracy, and enables reasoning by analogy to similar patients.
@@ Appendix A Additional Discussion & Limitations
There are limits to the number of experts that can be used in a Personalized MoE model for a given hidden dimension because: (i) the hidden dimension must be divisible by the number of experts, and (ii) the hidden dimension must be large enough to represent all discrete event times. Event times may cluster at key points in time, which would lower the dimensionality of the representation required to correctly decode. Our experiments work with 100 discrete event times and typically use a representation of size 16 (128 hidden dimensions / 8 experts), which is sufficient to represent the event time distribution for the datasets we explore, perhaps as a result of expert specialization by event time. If the event time distribution is more complex, then one may need number of dimensions on the order of number of time bins to form a sufficiently expressive event distribution. Another limitation is that we do not explore time-varying inputs and outputs (i.e., multiple predictions per patient as their input features are updated). In principle, all of our methods can be attached to a recurrent neural network or Transformer but the details of how to interpret groups of patients in a time-varying setting is an open question. Another pro/con of the Personalized MoE is that it is not as sensitive to the number of experts as the other two models, which is limiting if the primary goal is to discover latent groups. One heuristic is to use the Fixed MoE to discover the number of latent groups (e.g., using the elbow method shown in Figure 3) and then use that number of experts in the Personalized MoE model. Another approach is to use a large number of experts in the Personalized MoE model and then prune/consolidate experts based on their routing behavior, but we leave this for future work. In future work, we are also interested in exploring simpler (yet expressive) transformation families for the Adjustable MoE that are monotone and naturally map from $[0,1]$ to $[0,1]$.
@@ Appendix C Hyperparameters and Training Details
Table 5: Model specific hyperparameters for SUPPORT2 datasets.
@@
Table 6: Model specific hyperparameters for Sepsis datasets.

### v2

@@ Abstract
Deep mixture-of-experts models have attracted a lot of attention for survival analysis problems, particularly for their ability to cluster similar patients together. In practice, grouping often comes at the expense of key metrics such as calibration error and predictive accuracy. This is due to the restrictive inductive bias that mixture-of-experts imposes, that predictions for individual patients must look like predictions for the group they’re assigned to. Might we be able to discover patient group structure, where it exists, while improving calibration and predictive accuracy? In this work, we introduce several discrete-time deep mixture-of-experts (MoE)-based architectures for survival analysis problems, one of which achieves all desiderata: clustering, calibration, and predictive accuracy. We show that a key differentiator between this array of MoEs is how expressive their experts are. We find that more expressive experts that tailor predictions per patient outperform experts that rely on fixed group prototypes.
@@ 1 Introduction
The first architecture, Fixed MoE, uses several experts where each expert learns an associated event distribution, which is then fixed across all patients. The second architecture, Adjustable MoE, again learns a prototypical event distribution per expert, but can be adjusted to the individual patient. The third architecture, Personalized MoE, uses experts that each form a custom event distribution for individual patients. Only the third architecture is able to cluster patients and outperform strong baseline models on calibration, concordance index, and time-dependent Brier score. The contributions of our work are as follows: (i) we introduce three discrete-time deep MoE-based survival architectures, one of which achieves excellent clustering, calibration error, concordance index, and time-dependent Brier score and (ii) we report that the expressiveness of the experts is a key differentiator between discrete-time deep MoE-based survival analysis models, supported by a targeted set of experiments on these three architectures.
@@ 2 Related Work
Building survival models with flexible model classes, such as Gaussian processes (Alaa and van der Schaar, 2017) or neural networks (Ranganath et al., 2016), and rich distribution families (Miscouridou et al., 2018), yields better predictions of risk than standard linear Cox or accelerated failure-time (AFT) models. One can retain the structure of AFT models but improve prediction by using deep representations of the covariates (Zhong et al., 2021; Norman et al., 2024). Further generalizations that retain structure were introduced as conditional transformation models (Hothorn et al., 2018; Campanella et al., 2022). These choices improve expressivity to allow for better approximations of complicated survival distributions but do not reveal any structure in the data.
@@ 3 Methods
We now describe our three proposed deep mixture-of-experts (MoE)-based survival architectures. All models are feedforward deep learning models that use information from a patient’s record (e.g., demographic data, physiological data, etc.) to forecast when they are likely to have a clinical event. Raw patient records $\mathbf{x}_{0}^{i}\in\mathbb{R}^{a}$ for patients $i=1\ldots N$ and $a$ the number of raw features are composed of categorical indicators (e.g., gender, etc.), and standardized continuous features (e.g., heart rate), meaning that for continuous features we subtract the mean and divide by the standard deviation. We learn embeddings for categorical indicators. After embedding lookups are concatenated with continuous features, this results in a $d$-dimensional feature vector that is fed to the model. Let $\mathbf{x}=\mathbf{x}^{i}_{\ell-1}\in\mathbb{R}^{h}$ be shorthand for the $h$-dimensional penultimate hidden state representation of our feedforward model with $\ell$ layers for the $i^{\text{th}}$ patient. We will now describe the $\ell^{\text{th}}$ layer (i.e., the final layer), which is the MoE head in all architectures. All methods are trained using a discrete-time Multitask Logistic Regression (MTLR) style loss function (Yu et al., 2011; Fotso, 2018) that predicts a monotone label sequence (“once event occurs, it stays on”). This loss function has been shown to be well-calibrated (Haider et al., 2020). The loss functions are described in Appendix Section B.
@@ 3.3 Personalized Mixture-of-Experts
The Personalized MoE architecture belongs to a collection of models that provide the most flexibility to the experts to form custom event distributions per patient (e.g., Deep Survival Machines (Nagpal et al., 2021a)). Since this architecture generates new expert distributions for each patient, we first project the final hidden state representation $\mathbf{x}$ to a router representation with $\mathbf{x}_{r}=\mathbf{x}W^{\intercal}_{r}$ for $W_{r}\in\mathbb{R}^{h\times h}$ and an expert representation with $\mathbf{x}_{e}=\mathbf{x}W^{\intercal}_{e}$ for $W_{e}\in\mathbb{R}^{h\times h}$. We then divide the expert representation into $n$ evenly-sized chunks $\mathbf{x}_{e,k}$ for $k=1,\ldots,n$, which are each fed to a linear layer denoted $L_{k}\in\mathbb{R}^{m\times(h/n)}$ to form an event-distribution to obtain $M_{k}(\mathbf{x}_{e,k})=\mathbf{x}_{e,k}L_{k}^{\intercal}$, which collectively form the dynamic matrix of unnormalized densities over event times $M(\mathbf{x}_{e})\in\mathbb{R}^{n\times m}$. The final PMF is then
@@ 4 Experiments
We experiment with 3 datasets to probe the properties and capabilities of our proposed methods. Survival MNIST is a synthetic dataset (see Figure 2), which allows us to probe the models’ abilities to predict event times and recover latent groups. We censor 15% of examples. SUPPORT2 is a survival analysis dataset with 9,105 examples and $\sim$32% censoring (Connors et al., 1995). Sepsis is a larger dataset with 40,336 patient records and only 2,932 positive instances of sepsis, making this a very challenging anomaly detection survival analysis task (Reyna et al., 2020). For the Sepsis dataset, the first 100 hours of each patient’s ICU stay were summarized to perform a retrospective prediction, helpful when using partially labeled historical datasets to accelerate human labeling (Singer et al., 2016; Henry et al., 2019). Patients in the Sepsis dataset were administratively censored after 100 hours.
@@ 5 Results
We also analyze the clustering behavior of a Personalized MoE model trained on SUPPORT2 in Figure 5 to better understand how patient groups are formed on a real-world dataset. We see a variety of cluster sizes (Figure 5), indicating that Cluster 1 could possibly be merged into another cluster given its small size. We also find that patient risks are separated by cluster as evidenced by Figure 5. We then investigate kernel density estimates of continuous features (e.g., age in Figure 5) broken down by cluster. For categorical features, we define a $z$-score based on the proportion of patients in a cluster that have a given categorical feature value relative to the overall proportion of patients with that feature value (see Appendix Section J for more details). Importantly, these analyses allowed us to define succinct, clinician-informed signatures for each cluster, which are summarized in Table 2 and described in more detail in Appendix Section I. This indicates that the Personalized MoE model is able to discover clinically meaningful patient groups in real-world datasets.
@@ 6 Conclusion
We introduce three discrete-time deep mixture-of-experts (MoE)-based survival architectures, one of which achieves excellent clustering, calibration error, and Brier scores at key points in time. We have shown that the expressiveness of the experts is a key differentiator between deep MoE-based survival analysis models, supported by a targeted set of experiments on these three architectures. We observe that the Personalized MoE model is able to deliver on all desiderata: clustering, calibration, and predictive accuracy as evidenced by results on three survival analysis datasets, including two real-world datasets. We explore sensitivity to the number of experts and find that the Personalized MoE model is the least sensitive to the number of experts, making it a robust choice for survival analysis tasks. We also analyze the routing behavior of the MoE models and use a synthetic survival analysis dataset to validate that the models are able to recover latent groups in the data. In practice, we find that the Personalized MoE model is able to discover clinically meaningful patient groups in real-world datasets. Overall, the Personalized MoE model provides a powerful and flexible approach to survival analysis that can adapt to the complexities of real-world clinical data and patient heterogeneity. It maintains strong performance across multiple metrics of interest to clinicians and researchers alike, including calibration and predictive accuracy, and enables reasoning by analogy to similar patients.
@@ Appendix A Additional Discussion & Limitations
There are limits to the number of experts that can be used in a Personalized MoE model for a given hidden dimension because: (i) the hidden dimension must be divisible by the number of experts, and (ii) the hidden dimension must be large enough to represent all discrete event times. Event times may cluster at key points in time, which would lower the dimensionality of the representation required to correctly decode. Our experiments work with 100 discrete event times and typically use a representation of size 16 (128 hidden dimensions / 8 experts), which is sufficient to represent the event time distribution for the datasets we explore, perhaps as a result of expert specialization by event time. If the event time distribution is more complex, then one may need a number of dimensions on the order of number of time bins to form a sufficiently expressive event distribution. Another limitation is that we do not explore time-varying inputs and outputs (i.e., multiple predictions per patient as their input features are updated). In principle, all of our methods can be attached to a recurrent neural network or Transformer but the details of how to interpret groups of patients in a time-varying setting is an open question. Another pro/con of the Personalized MoE is that it is not as sensitive to the number of experts as the other two models, which is limiting if the primary goal is to discover latent groups. One heuristic is to use the Fixed MoE to discover the number of latent groups (e.g., using the elbow method shown in Figure 3) and then use that number of experts in the Personalized MoE model. Another approach is to use a large number of experts in the Personalized MoE model and then prune/consolidate experts based on their routing behavior, but we leave this for future work. In future work, we are also interested in exploring simpler (yet expressive) transformation families for the Adjustable MoE that are monotone and naturally map from $[0,1]$ to $[0,1]$.
@@ Appendix C Hyperparameters and Training Details
Table 5: Model specific hyperparameters for SUPPORT2 dataset.
@@
Table 6: Model specific hyperparameters for Sepsis dataset.